If 10g of calcium chloride us dissolved in 300g of water to make solution with density 1.2g/cc then it's molarity will be?

Grade:11

1 Answers

Sahani Kumar
5 years ago
Mass of the CaCl2 = 10gram
Molecular mass of cacl2=40 + (2*(35.5))=111
Vant Hoff factor for CaCl2 = 3 (because CaCl2.                                                     dissociate into 3 particle.)
Now, 
Number of moles
of CaCl2 = (mass of CaCl2) /
                                 (  Molecular mass of CaCl2)
So number of moles if CaCl2 = 10/111=0.09
Volume of water = 300g which is = 300ml or 0.3L
Now finally
Molarity =[ (Vant Hoff factor)*(number of moles)]/                                                   [Volume]
Molarity=  [(3)*(0.09)]/[0.3] = 0.9
Molarity = 0.9
Regards 
Askiitian member
If you liked this answer please encourage us by approving this answer.